# Descending finite projective modules from a Novikov ring

### v1

@@ 4.4. Restricting the exponents
we see that $\operatorname{supp}(x)$ must be contained in ×$\mathbb{R}\times\Gamma$. For every element $m\in M$, we in particular have $\operatorname{supp}(\pi_{2}^{\ast}m)\subseteq\mathbb{R}\times\Gamma$, and hence $\operatorname{supp}(m)\subseteq\Gamma$. This shows that
@@ 5. Application to perfectoid geometry
Let $A$ be a perfect $\mathbb{F}_{p}$-algebra. Then the natural functor
{vector bundles on $\{\text{vector bundles on }\operatorname{Spec}A\}\to\{\text{vector bundles on }\operatorname{Spd}A\}$
@@
is a v-cover, by [SW20, Lemma 17.1.8] the category of vector bundles on $\operatorname{Spd}A$ can be identified with the category of pairs $(\mathscr{V},\varphi)$, where $\mathscr{V}$ is a vector bundle on $X$ and $\varphi$ is an isomorphism of vector bundles over $X\times_{Y}X$ satisfying a cocycle condition over $X\times_{Y}X\times_{Y}X$. By [KL15, Theorem 2.7.7] and again [SW20, Lemma 17.1.8], the vector bundle $\mathscr{V}$ corresponds to a finite projective module $M$ over $A(\!(t^{{1/p^{\infty}}})\!)$.
Next, we identify
$X\times_{Y}X=\bigcup_{n=2}^{\infty}\operatorname{Spd}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\Bigl{\langle}\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr{\rangle}[1/tu].$
If we denote by $\pi_{1},\pi_{2}\colon X\times_{Y}X\to X$ the two projection maps, then $\varphi$ is an isomorphism of vector bundles
$\varphi\colon\pi_{1}^{\ast}\mathscr{V}\xrightarrow{\cong}\pi_{2}^{\ast}\mathscr{V}$
over $X\times_{Y}X$. Because
$H^{0}(X\times_{Y}X,\mathscr{O}_{X\times_{Y}X})=\bigcap_{n=2}^{\infty}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\Bigl{\langle}\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr{\rangle}[1/tu]=A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!),$
we see that $\varphi$ corresponds to an isomorphism
@@
Finally the cocycle condition is defined on
$X\times_{Y}X\times_{Y}X=\bigcup_{n=3}^{\infty}\operatorname{Spd}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}},v^{{1/p^{\infty}}}]]\Bigl{\langle}\frac{t^{n}}{uv},\frac{u^{n}}{tv},\frac{v^{n}}{tu}\Bigr{\rangle}[1/tuv],$
and hence be checked after tensoring to
@@
This shows that a vector bundle on $\operatorname{Spd}A$ corresponds to a Novikov descent datum for $A$. Applying Theorem 4.9 now gives the desired result. ∎

### v2

@@ 4.4. Restricting the exponents
we see that $\operatorname{supp}(x)$ must be contained in $\mathbb{R}\times\Gamma$. For every element $m\in M$, we in particular have $\operatorname{supp}(\pi_{2}^{\ast}m)\subseteq\mathbb{R}\times\Gamma$, and hence $\operatorname{supp}(m)\subseteq\Gamma$. This shows that
@@ 5. Application to perfectoid geometry
Let $A$ be a discrete perfect $\mathbb{F}_{p}$-algebra. There is a v-cover
$X=\operatorname{Spd}(A(\!(t^{{1/p^{\infty}}})\!),A[[t^{{1/p^{\infty}}}]])\to\operatorname{Spd}(A,A)=Y$
of $Y$ be an affinoid perfectoid space $X$. We note that $X$ represents the functor sending a perfectoid Huber pair $(A,A)\to(R,R^{+})$ to a topologically nilpotent unit $\varpi\in R^{\circ\circ}$, because uniformity and perfectness of $R$ implies that there exists a unique continuous map $A(\!(t^{{1/p^{\infty}}})\!)\to R$ sending $t\mapsto\varpi$. This means that both $X\times_{Y}X$ and $X\times_{Y}X\times_{Y}X$ are representable by perfectoid spaces
$X\times_{Y}X$ $=\operatorname{Spa}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\setminus\{\lvert tu\rvert=0\}$
$=\bigcup_{n=2}^{\infty}\operatorname{Spd}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\Bigl\langle\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr\rangle[1/tu],$
$X\times_{Y}X\times_{Y}X$ $=\operatorname{Spa}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}},v^{{1/p^{\infty}}}]]\setminus\{\lvert tuv\rvert=0\}$
$=\bigcup_{n=3}^{\infty}\operatorname{Spd}A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}},v^{{1/p^{\infty}}}]]\Bigl\langle\frac{t^{n}}{uv},\frac{u^{n}}{tv},\frac{v^{n}}{tu}\Bigr\rangle[1/tuv].$
Lemma 5.2.
We have
$H^{0}(X\times_{Y}X,\mathscr{O}_{X\times_{Y}X})$ $=A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!),$
Proof.
For the first equality, we note that we can write
$A[[t^{{1/p^{\infty}}},u^{{1/p^{\infty}}}]]\Bigl\langle\frac{t^{n}}{u},\frac{u^{n}}{t}\Bigr\rangle\Bigl[\frac{1}{tu}\Bigr]=\left\{\sum_{i\in I}a_{i}t^{d_{i}}u^{e_{i}}:\begin{matrix}a_{i}\in A,d_{i},e_{i}\in\mathbb{Z}[p^{-1}],\\ \{i\in I:\alpha d_{i}+e_{i}<C\}\text{ is finite}\\ \text{for all }C\in\mathbb{R}\text{ and }n^{-1}\leq\alpha\leq n\end{matrix}\right\}.$
As we increase the $n$ and take the intersection of the rings, we see that we allow all possible values of $\alpha>0$, and hence the ring of global sections of $X\times_{Y}X$ is $A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!)$. The proof of the second equality is similar. ∎
@@
Let $A$ be a discrete perfect $\mathbb{F}_{p}$-algebra. Then the natural functor
$\{\text{vector bundles on }\operatorname{Spec}A\}\to\{\text{vector bundles on }\operatorname{Spd}A\}$
@@
from above. By [SW20, Lemma 17.1.8] the category of vector bundles on $\operatorname{Spd}A$ can be identified with the category of pairs $(\mathscr{V},\varphi)$, where $\mathscr{V}$ is a vector bundle on $X$ and $\varphi$ is an isomorphism of vector bundles over $X\times_{Y}X$ satisfying a cocycle condition over $X\times_{Y}X\times_{Y}X$. By [KL15, Theorem 2.7.7] and again [SW20, Lemma 17.1.8], the vector bundle $\mathscr{V}$ corresponds to a finite projective module $M$ over $A(\!(t^{{1/p^{\infty}}})\!)$.
By Lemma 5.2, the natural functor from the category of finite projective $A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!)$-modules to the category of vector bundles on $X\times_{Y}X$ is fully faithful. If we denote by $\pi_{1},\pi_{2}\colon X\times_{Y}X\to X$ the two projection maps, then the descent datum $\varphi\colon\pi_{1}^{\ast}\mathscr{V}\cong\pi_{2}^{\ast}\mathscr{V}$ corresponds to an isomorphism of $A(\!(t^{{1/p^{\infty}}},u^{{1/p^{\infty}}})\!)$-modules
@@
Finally the cocycle condition is defined on $X\times_{Y}X\times_{Y}X$, and hence by Lemma 5.2 can be checked after tensoring to
@@
This shows that a vector bundle on $\operatorname{Spd}A$ corresponds to a Novikov descent datum for $A$ with $\Gamma=\mathbb{Z}[p^{-1}]$. Applying Theorem 1.1 now gives the desired result. ∎